# Parafermions with symmetry-protected non-Abelian statistics

### 14 Mar 2024 v1

@@ Abstract
Non-Abelian anyons have garnered extensive attention for obeying exotic non-Abelian statistics and potential applications to fault-tolerant quantum computation. Although the prior research has predominantly focused on non-Abelian statistics without the necessity of symmetry protection, recent progresses have shown that symmetries can play essential roles and bring about a notion of the symmetry-protected non-Abelian (SPNA) statistics. In this work, we extend the concept of SPNA statistics to strongly-correlated systems which host parafermion zero modes (PZMs). This study involves a few fundamental results proved here. First, we unveil a generic unitary symmetry mechanism that protects PZMs from local couplings. Then, with this symmetry protection, the PZMs can be categorized into two nontrivial sectors, each maintaining its own parity conservation, even though the whole system cannot be dismantled into separate subsystems due to nonlinear interactions. Finally, by leveraging the parity conservation of each sector and the general properties of the effective braiding Hamiltonian, we prove rigorously that the PZMs intrinsically obey SPNA statistics. To further confirm the results, we derive the braiding matrix at a tri-junction. In addition, we propose a physical model that accommodates a pair of PZMs protected by mirror symmetry and satisfying the generic theory. This work shows a broad spectrum of strongly-correlated systems capable of hosting fractional SPNA quasiparticles and enriches our comprehension of fundamental quantum statistics linked to the symmetries that govern the exchange dynamics.
@@ I Introduction
Among various non-Abelian anyons, the Majorana zero modes (MZMs) stand out due to the recent considerable experimental progress in the topological superconductors Mourik2012 ; MTDeng2012 ; Rokhinson2012 ; Shtrikman2012 ; JFJia2012 ; Marcus2013 ; JFJia2014 ; Yazdani2014 ; Marcus2015 ; Marcus2016 ; Molenkamp2016 ; Molenkamp2017 ; Qiao2017 ; HDing2018a ; HDing2018b ; Marcus2019 ; Molenkamp2019 ; Yazdani2019 ; Zhang2022 ; Microsoft2023 , where the Majorana modes emerge as quasiparticles identical to their antiparticles Kitaev2001 ; AliceaRPP ; Flensberg2012 ; Beenakker2013 ; Franz2015 ; Sato2017 . Two MZMs define a non-local complex fermion. The qubits encoded by these non-local fermions are robust against local perturbations Kitaev2001 . Braiding and measuring MZMs can implement Clifford gates, although these are not sufficient for universal quantum computation Gottesman1998 . From an algebraic standpoint, MZMs are considered as a special case within the broader category of $\mathbb{Z}_{N}$ parafermion zero modes (PZMs) with $N=2$ FradkinKadanoff ; Fendley2012 . Physically, the parafermions are fractionalized MZMs, emerging only in the strongly correlated systems SternPRX2012 ; MCheng2012 ; AliceaNC2013 ; Barkeshli2013 ; Barkeshli2014 ; Stern2014 ; Loss2014a ; Loss2014b ; Loss2014c ; Oreg2017general ; Oreg2017chiral ; Loss2019 ; Santos2020 ; Loss2020bilayer ; Oreg2020predicted ; Silva2022 . The braiding phases of PZMs also exhibit fractionalization compared to those of MZMs, manifesting a kind of non-Abelian fractional statistics. The parafermions can encode the nonlocal topological qudits FockPara , and braiding parafermion brings us closer to the computational universality than braiding MZMs AliceaNC2013 ; Loss2016 .
@@
The non-Abelian statistics was initially introduced to singly existing MZMs which necessitate no symmetry protection Ivanov2001 ; Alicea2011 . The presence of symmetries can enrich the broad classes of topological superconductors SCZhang2009 ; TeoKane2010 ; Timm2010 ; Beenakker2011 ; KTLaw2012 ; Nagaosa2012 ; KaneMele2013 ; Berg2013 ; DLoss-a ; DLoss-b ; Nagaosa2014 ; magneticTSC ; Oreg2019 ; mirrorTSC ; Ueno2013 ; Sato2014crystalline ; C4TSC ; InversionTSC , in which a single topological defect may bind multiple zero modes and generate the ground state degeneracy. It was first shown in the time-reversal invariant topological superconductors that the Majorana Kramers pairs obey non-Abelian statistics due to the protection of time-reversal symmetry, rendering a new family of quantum statistics called the symmetry-protected non-Abelian (SPNA) statistics XJL2014PRX ; Gao2016 . This prediction shows an insight into the nontrivial roles of symmetries in the quantum statistics. The SPNA statistics has been further considered in various symmetry-protected topological superconductors Sato2014mirror ; JSH2022 ; YWuReview2023 ; NittaReview2024 , and is also the essential mechanism JSH2022 underlying the braiding statistics of Dirac zero modes Nitta2012dirac ; Loss2013fractional ; YWu2020a ; YWu2020b ; YWu2022 . The unitary and anti-unitary symmetries are fundamentally different in protecting the non-Abelian statistics. For the anti-unitary symmetric system, e.g. the time-reversal invariant topological superconductor, the non-Abelian statistics of Majorana Kramers pairs requires that not only the static topological superconductor satisfy the time-reversal symmetry, but also the braiding operation, which characterizes a dynamical evolution, satisfy time-reversal symmetry individually Gao2016 . This is because the time-reversal symmetry for the static superconductor cannot guarantee the absence of local coupling between two MZMs of each Kramers pair in the dynamical evolution localmixing1 ; localmixing2 ; Knapp2020 due to the profound dynamical symmetry-breaking of anti-unitary symmetries Gao2016 ; JasonAlicea2020. The dynamical symmetry breaking with coupling to environment opens up active discussions on the robustness of time-reversal invariant topological phases in the open systems Cooper2020 ; Zhai2021 ; Cai2021 . In contrast, the Majorana modes protected by unitary symmetries are intrinsic SPNA anyons without suffering the dynamical symmetry breaking JSH2022 . In particular, when the unitary symmetry is non-Abelian, the whole system cannot be block diagonalized into decoupled sectors corresponding to different symmetry eigenvalues, while the non-Abelian statistics is generically protected by the symmetry JSH2022 . The SPNA statistics elucidates that the quantum statistics of indistinguishable (quasi)-particles is essentially a dynamical rather than static fundamental feature. In additional to fundamental relevance, the symmetry-protected MZMs enrich new schemes for topological quantum computing JSH2022 ; OregReview2019 ; MCheng2021 ; TanakaMKP2022 ; LFu2022 .
@@ II Non-Abelian braiding of PZMs with unitary symmetry protection
In this section, through a systematic study we prove rigorously that PZMs with unitary symmetry protection are intrinsic non-Abelian fractional quasiparticles. We first unveil the generic mechanism of unitary symmetry protection for the locally coexisting PZMs, for which all the local couplings are forbidden. This result leads to two key consequences. One is that the PZMs can be categorized into two separate sectors by redefining these modes from a nontrivial symmetric decomposition, yet the whole system cannot be decoupled into independent subsystems due to the strong interactions. The other is that in terms of the decomposed PZMs, the anyon parity of each symmetry sector is conserved. With the parity conservation and topological spin statistics theorem, we prove rigorously that the PZMs with unitary symmetry protection intrinsically obey the SPNA statistics.
@@ II.1 Unitary symmetry protection
We first define the system of locally-coexisting PZMs and determine the generic form of the local coupling terms without symmetry protection. Consider two pairs of $\mathbb{Z}_{N}$ PZMs localized at two edges (domain walls) of the system with the commutation relations
@@
where $\omega=\exp(2\pi i/N)$, $\tilde{\alpha}_{i=1,2}$, $\tilde{\beta}_{j=1,2}$ are unitary operators that commute with the system’s Hamiltonian and satisfy $\tilde{\alpha}_{i}^{N}=\tilde{\beta}_{i}^{N}=1$, and $\tilde{\alpha}_{1(2)}$, $\tilde{\beta}_{1(2)}$ localize at the left (right) edge of the system. Without symmetry protection, the PZMs $\tilde{\alpha}$ and $\tilde{\beta}$ of different flavours may be coupled together by local Hamiltonian $H_{\text{c}}=\sum_{k,l}\xi_{k,l}\tilde{\alpha}_{1}^{k}\tilde{\beta}_{1}^{l}+\sum_{m,n}\eta_{m,n}\tilde{\alpha}_{2}^{m}\tilde{\beta}_{2}^{n}$, where $\xi_{k,l}^{*}=\omega^{kl}\xi_{N-k,N-l}$ and $\eta_{m,n}^{*}=\omega^{mn}\eta_{N-m,N-n}$ for hermitianity. Since the coupling terms must conserve the total parity defined by $\tilde{Q}_{\text{tot}}=\omega\tilde{\alpha}_{1}\tilde{\alpha}_{2}^{\dagger}\tilde{\beta}_{1}\tilde{\beta}_{2}^{\dagger}$, we can determine generally the concrete form of the local couplings
@@
where $\xi_{m}=\xi_{N-m}^{*}$ and $\eta_{n}=\eta_{N-n}^{*}$ for hermitianity and $\tilde{Q}_{i}=\omega^{(N+1)/2}\tilde{\alpha}_{i}\tilde{\beta}_{i}^{\dagger}$ denotes the local $\mathbb{Z}_{N}$ parity. The Hamiltonian Eq. (3) describes the generic local coupling terms in the absence of symmetry protection other than the intrinsic total parity conservation. The presence of the these terms pushes the parafermion modes out of zero energy, and the role of symmetry protection is then to forbid these terms as detailed later.
We show that the generic unitary symmetry protection condition for a pair of PZMs can be expressed by the action of a unitary symmetry at the local parity operator. By definition, the most general symmetry protection of these PZMs is $[H_{\text{c}},S]\neq 0$, for $H_{\text{c}}$ in any form, where $S$ is a symmetry of the system. When $N=2$, $\tilde{\alpha}_{i}$ and $\tilde{\beta}_{i}$ are reduced to MZMs, then only two terms $i\xi_{1}\tilde{\alpha}_{1}\tilde{\beta}_{1}$ and $i\eta_{1}\tilde{\alpha}_{2}\tilde{\beta}_{2}$ survive, and they can be eliminated by a unitary symmetry $Si\tilde{\alpha}_{i}\tilde{\beta}_{i}S^{-1}=-i\tilde{\alpha}_{i}\tilde{\beta}_{i}$ JSH2022 . When $N>2$, there is more than one coupling term at a single edge, thus the symmetry protection condition is more complicate. To facilitate the derivation, we work in the bases of eigenstates of local parity operator $\tilde{Q}_{i}$. The generic condition $[H_{\text{c}},S]\neq 0$ implies that the unitary symmetry $S$ must be a cyclic permutation of the eigenstates with $N$ elements provided that $S|q\rangle$ is still a local parity eigenstate. In this case, there exists a unitary matrix $G$ such that $GSG^{-1}=e^{i\vartheta/N}V^{p}$, where $p$ and $N$ are coprime, $\vartheta$ is a phase factor and $V$ is the shift matrix $V|q\rangle=|q-1\rangle$ (the details are provided in Appendix A). This observation can be rewritten in the operator form
@@
where $Q_{i}=G^{-1}\tilde{Q}_{i}G$ is the local parity operator of PZMs $\alpha^{\prime}_{i}=G^{-1}\tilde{\alpha}_{i}G$ and $\beta^{\prime}_{i}=G^{-1}\tilde{\beta}_{i}G$. This condition Eq. (4) suggests that the symmetry can advance the new local parity $Q_{i}$ by $p$-increments, where $p$ and $N$ are coprime. For $N=2$, this means that the symmetry operation switches odd/even fermion parity to even/odd one.
II.2 Parity conservation for each sector
In this subsection, we prove that the generic condition for unitary symmetry protection Eq. (4) implies that the PZMs can be classified into two sectors, with the parity of each sector being conserved. We can show generically that the unitary symmetry $S$ operates on the PZMs as (see Appendix B for more details)
@@ II.2 Parity conservation for each sector
where $\mu_{n}^{*}=\mu_{N-n}$, $p_{1}$, $p_{2}$ are integers that satisfy $p_{1}-p_{2}=p$, and $\alpha^{\prime}_{i}=G^{-1}\tilde{\alpha}_{i}G$, $\beta^{\prime}_{i}=G^{-1}\tilde{\beta}_{i}G$. We see that the symmetry operator couples $\alpha^{\prime}_{i}$ and $\beta^{\prime}_{i}$. By defining new PZMs via unitary transformation $\alpha_{i}=W\alpha^{\prime}_{i}W^{-1}$, $\beta_{i}=W\beta^{\prime}_{i}W^{-1}$, where $W=\exp\left({i\sum_{n=1}^{N-1}\nu_{n}Q_{i}^{n}}\right)$, $\nu_{n}=-\mu_{n}/[(\omega^{n}-1)(\omega^{pn}-1)]$, we find the symmetry protection condition can be rewritten in a decoupled form
$S\alpha_{i}S^{-1}=\omega^{p_{1}}\alpha_{i},\quad S\beta_{i}S^{-1}=\omega^{p_{2}}\beta_{i},$ (7)
which indicates that $\alpha_{i}$ and $\beta_{i}$ are eigenmodes of the symmetry. When $N=2$, the above procedure can be reduced to the Majorana case. For a pair of MZMs $\alpha^{\prime}_{i}$ and $\beta^{\prime}_{i}$, the symmetry acts as $S\alpha^{\prime}_{i}S^{-1}=\pm(\cos\phi\alpha^{\prime}_{i}+\sin\phi\beta^{\prime}_{i})$, $S\beta^{\prime}_{i}S^{-1}=\mp(-\sin\phi\alpha^{\prime}_{i}+\cos\phi\beta^{\prime}_{i})$, which give $\mu_{1}=\phi$. Then by setting $\nu_{1}=-\phi/4$, we have $\alpha_{i}=\cos\frac{\phi}{2}\alpha^{\prime}_{i}+\sin\frac{\phi}{2}\beta^{\prime}_{i}$, $\beta_{i}=-\sin\frac{\phi}{2}\alpha^{\prime}_{i}+\cos\frac{\phi}{2}\beta^{\prime}_{i}$, and it’s easy to verify that $S\alpha_{i}S^{-1}=\pm\alpha_{i}$, $S\beta_{i}S^{-1}=\mp\beta_{i}$. In other words, the MZMs $\alpha_{i}$ and $\beta_{i}$ are eigenstates corresponding to different eigenvalues ±$\pm 1$ of the unitary symmetry.
According to the decoupled form of unitary symmetry protection Eq. (7), the parity of each sector defined by $Q_{\alpha}=\omega^{(N+1)/2}\alpha^{\prime}_{1}{\alpha^{\prime}}_{2}^{\dagger}=\omega^{(N+1)/2}\alpha_{1}\alpha_{2}^{\dagger}$, $Q_{\beta}=\omega^{(N+1)/2}\beta^{\prime}_{1}{\beta^{\prime}}_{2}^{\dagger}=\omega^{(N+1)/2}\beta_{1}\beta_{2}^{\dagger}$ should be conserved. To see this, consider the generic non-local four-body interaction term $H_{I}=\xi_{k,l,m,n}\alpha_{1}^{k}\alpha_{2}^{l}\beta_{1}^{m}\beta_{2}^{n}+h.c.$, where $k+l+m+n=0\text{ mod }N$ due to the total parity conservation. The symmetry protection condition Eq. (7) further requires $k+l=m+n=0\text{ mod }N$, then the interaction term reads $H_{I}=\xi_{k,l,m,n}Q_{\alpha}^{k}Q_{\beta}^{m}+h.c.$, which commutes with the parity operator of each sector. Note that the interaction term couples $\alpha_{i}$-sector and $\beta_{i}$-sector when $k\neq 0$, $m\neq 0$, we remark that the parity conservation of each sector does not imply the decomposition of the whole system into two distinct subsystems.

### 21 Mar 2024 v2

@@ Abstract
Non-Abelian anyons have garnered extensive attention for obeying exotic non-Abelian statistics and having potential applications to fault-tolerant quantum computing. While the prior research has predominantly focused on non-Abelian statistics without the necessity of symmetry protection, recent progresses have shown that symmetries can play essential roles and bring a notion of the symmetry-protected non-Abelian (SPNA) statistics. In this work, we extend the concept of SPNA statistics to strongly-correlated systems which host parafermion zero modes (PZMs). This study involves a few fundamental results proved here. First, we unveil a generic unitary symmetry mechanism that protects PZMs from local couplings. Then, with this symmetry protection, the PZMs can be categorized into two nontrivial sectors, each maintaining its own parity conservation, even though the whole system cannot be dismantled into separate subsystems due to nonlinear interactions. Finally, by leveraging the parity conservation of each sector and the general properties of the effective braiding Hamiltonian, we prove rigorously that the PZMs intrinsically obey SPNA statistics. To further confirm the results, we derive the braiding matrix at a tri-junction. We also propose a correlated quantum nanowire model that accommodates a pair of PZMs protected by mirror symmetry and satisfying the generic theory. This work shows a broad spectrum of strongly-correlated systems capable of hosting fractional SPNA quasiparticles and enriches our comprehension of fundamental quantum statistics linked to the symmetries that govern the exchange dynamics.
@@ I Introduction
Among various non-Abelian anyons, the Majorana zero modes (MZMs) stand out due to the recent considerable experimental progress in the topological superconductors Mourik2012 ; MTDeng2012 ; Rokhinson2012 ; Shtrikman2012 ; JFJia2012 ; Marcus2013 ; JFJia2014 ; Yazdani2014 ; Marcus2015 ; Marcus2016 ; Molenkamp2016 ; Molenkamp2017 ; HDing2018a ; HDing2018b ; Marcus2019 ; Molenkamp2019 ; Yazdani2019 ; Zhang2022 ; Microsoft2023 , where the Majorana modes emerge as quasiparticles identical to their antiparticles Kitaev2001 ; AliceaRPP ; Flensberg2012 ; Beenakker2013 ; Franz2015 ; Sato2017 ; YPHe2020 . Two MZMs define a non-local complex fermion. The qubits encoded by these non-local fermions are robust against local perturbations Kitaev2001 . Braiding and measuring MZMs can implement Clifford gates, although these are not sufficient for universal quantum computation Gottesman1998 . From an algebraic standpoint, MZMs are considered as a special case within the broader category of $\mathbb{Z}_{N}$ parafermion zero modes (PZMs) with $N=2$ FradkinKadanoff ; Fendley2012 ; Motruk2013topological ; Bondesan2013 ; Meidan2017 ; GMZhang2018a . Physically, the parafermions are fractionalized MZMs, emerging only in the strongly correlated systems SternPRX2012 ; MCheng2012 ; AliceaNC2013 ; Barkeshli2013 ; Barkeshli2014 ; Stern2014 ; Loss2014a ; Loss2014b ; Loss2014c ; Oreg2017general ; Oreg2017chiral ; Loss2019 ; Santos2020 ; Loss2020bilayer ; Oreg2020predicted ; Silva2022 ; Gefen2022 . The braiding phases of PZMs also exhibit fractionalization compared to those of MZMs, manifesting a kind of non-Abelian fractional statistics. The parafermions can encode the nonlocal topological qudits FockPara , and braiding parafermion brings us closer to the computational universality than braiding MZMs AliceaNC2013 ; Loss2016 .
@@
The non-Abelian statistics was initially introduced to singly existing MZMs which necessitate no symmetry protection Nayak1996 ; Ivanov2001 ; Alicea2011 . The presence of symmetries can enrich the broad classes of topological superconductors SCZhang2009 ; TeoKane2010 ; Timm2010 ; Beenakker2011 ; KTLaw2012 ; Nagaosa2012 ; KaneMele2013 ; Berg2013 ; DLoss-a ; DLoss-b ; Nagaosa2014 ; magneticTSC ; Oreg2019 ; mirrorTSC ; Ueno2013 ; Sato2014crystalline ; C4TSC ; InversionTSC , in which a single topological defect may bind multiple zero modes and generate the ground state degeneracy. It was first shown in the time-reversal invariant topological superconductors that the Majorana Kramers pairs obey non-Abelian statistics due to the protection of time-reversal symmetry, rendering a new family of quantum statistics called the symmetry-protected non-Abelian (SPNA) statistics XJL2014PRX ; Gao2016 . This prediction shows an insight into the nontrivial roles of symmetries in the quantum statistics. The SPNA statistics has been further considered in various symmetry-protected topological superconductors Sato2014mirror ; JSH2022 ; YWuReview2023 ; NittaReview2024 , and is also the essential mechanism JSH2022 underlying the braiding statistics of Dirac zero modes Nitta2012dirac ; Loss2013fractional ; YWu2020a ; YWu2020b ; YWu2022 . The unitary and anti-unitary symmetries are fundamentally different in protecting the non-Abelian statistics. For the system with anti-unitary symmetry, e.g. the time-reversal invariant topological superconductor, the non-Abelian statistics of Majorana Kramers pairs requires that not only the static topological superconductor satisfy the time-reversal symmetry, but also the braiding operation, which characterizes a dynamical evolution, satisfy a time-reversal symmetry individually Gao2016 . This is because the time-reversal symmetry for the static superconductor cannot exclude the local coupling between two MZMs of a single Kramers pair in the generic dynamical evolution localmixing1 ; localmixing2 ; Knapp2020 due to the profound dynamical symmetry-breaking of anti-unitary symmetries Gao2016 ; JasonAlicea2020 . The dynamical symmetry breaking with coupling to environment opens up an interesting issue on the robustness of time-reversal invariant topological phases in the open systems Cooper2020 ; Zhai2021 ; Cai2021 . In contrast, the Majorana modes protected by unitary symmetries are intrinsic SPNA anyons without suffering the dynamical symmetry breaking. In particular, when the unitary symmetry is non-Abelian, the entire system cannot be block diagonalized into decoupled sectors corresponding to different symmetry eigenvalues, but the non-Abelian statistics is fully protected by the symmetry JSH2022 . The SPNA statistics elucidates that the quantum statistics of indistinguishable (quasi)-particles is essentially a dynamical rather than a static feature. In additional to the fundamental relevance, the symmetry-protected MZMs enrich new schemes for topological quantum computing JSH2022 ; OregReview2019 ; MCheng2021 ; TanakaMKP2022 ; LFu2022 .
@@ II Non-Abelian braiding of PZMs with unitary symmetry protection
In this section, through a systematic study we prove rigorously that PZMs with unitary symmetry protection are intrinsic non-Abelian fractional quasiparticles. We first unveil the generic mechanism of unitary symmetry protection for the locally coexisting PZMs, for which all the local couplings are forbidden. This result leads to two key consequences. One is that the PZMs can be categorized into two separate sectors by redefining these modes from a sophisticated symmetric decomposition, yet the whole system cannot be decoupled into independent subsystems due to the strong interactions. The other is that in terms of the decomposed PZMs, the anyon parity of each symmetry sector is conserved. With the parity conservation and topological spin statistics theorem, we prove rigorously that the PZMs with unitary symmetry protection intrinsically obey the SPNA statistics.
@@ II.1 Unitary symmetry protection
We first introduce the system of locally-coexisting PZMs and determine the generic form of the local coupling terms breaking symmetry protection. Consider two pairs of $\mathbb{Z}_{N}$ PZMs localized at two edges (domain walls) of the system with the commutation relations
@@
where $\omega=\exp(2\pi i/N)$, $\tilde{\alpha}_{i=1,2}$, $\tilde{\beta}_{j=1,2}$ are unitary operators that commute with the system’s Hamiltonian and satisfy $\tilde{\alpha}_{i}^{N}=\tilde{\beta}_{i}^{N}=1$, and $\tilde{\alpha}_{1(2)}$, $\tilde{\beta}_{1(2)}$ localize at the left (right) edge of the system. Without symmetry protection, the PZMs $\tilde{\alpha}$ and $\tilde{\beta}$ of different flavours may be coupled together by local Hamiltonian $H_{\text{c}}=\sum_{k,l}\xi_{k,l}\tilde{\alpha}_{1}^{k}\tilde{\beta}_{1}^{l}+\sum_{m,n}\eta_{m,n}\tilde{\alpha}_{2}^{m}\tilde{\beta}_{2}^{n}$, where $\xi_{k,l}^{*}=\omega^{kl}\xi_{N-k,N-l}$ and $\eta_{m,n}^{*}=\omega^{mn}\eta_{N-m,N-n}$ for hermitianity. Since the coupling terms must conserve the total parity defined by $\tilde{Q}_{\text{tot}}=\omega\tilde{\alpha}_{1}\tilde{\alpha}_{2}^{\dagger}\tilde{\beta}_{1}\tilde{\beta}_{2}^{\dagger}$, we can write down the concrete form of the local couplings as
@@
where $\xi_{m}=\xi_{N-m}^{*}$ and $\eta_{n}=\eta_{N-n}^{*}$ for hermitianity and $\tilde{Q}_{i}=\omega^{(N+1)/2}\tilde{\alpha}_{i}\tilde{\beta}_{i}^{\dagger}$ denotes the local $\mathbb{Z}_{N}$ parity. The Hamiltonian Eq. (3) describes the generic local coupling terms in the absence of symmetry protection other than the intrinsic total parity conservation. The presence of the these terms pushes the parafermion modes out of zero energy, and the role of symmetry protection is then to forbid these terms as detailed below.
We show that the generic unitary symmetry protection for any pair of PZMs can be expressed through the unitary transformation on the local anyon parities. If there is a unitary symmetry $S$ protecting the PZMs, it implies that the local coupling of these modes are prohibited by requiring that $[H_{\text{c}},S]\neq 0$ for arbitrary $H_{\text{c}}$. When $N=2$, $\tilde{\alpha}_{i}$ and $\tilde{\beta}_{i}$ are reduced to MZMs. Then $H_{c}$ only involves the terms $i\xi_{1}\tilde{\alpha}_{1}\tilde{\beta}_{1}$ and $i\eta_{1}\tilde{\alpha}_{2}\tilde{\beta}_{2}$, which can be eliminated by introducing a unitary symmetry $Si\tilde{\alpha}_{i}\tilde{\beta}_{i}S^{-1}=-i\tilde{\alpha}_{i}\tilde{\beta}_{i}$. For PZMs with $N>2$, the symmetry protection mechanism is generally more complicate. To facilitate the derivation, we work in the bases of eigenstates $|q\rangle$ of local parity operator $\tilde{Q}_{i}$. The generic condition $[H_{\text{c}},S]\neq 0$ implies that the unitary symmetry $S$ must be a cyclic permutation of the eigenstates with $N$ elements provided that $S|q\rangle$ is still a local parity eigenstate. In this case, there exists a unitary matrix $G$ such that $GSG^{-1}=e^{i\vartheta/N}V^{p}$, where $p$ and $N$ are coprime, $\vartheta$ is a phase factor and $V$ is the shift matrix $V|q\rangle=|q-1\rangle$ (the details are provided in Appendix A). This observation can finally be rewritten in the form
@@
where $Q_{i}=G^{-1}\tilde{Q}_{i}G$ is the local parity operator of PZMs $\alpha^{\prime}_{i}=G^{-1}\tilde{\alpha}_{i}G$ and $\beta^{\prime}_{i}=G^{-1}\tilde{\beta}_{i}G$. The Eq. (4) suggests that the symmetry can advance the new local parity $Q_{i}$ by $p$-increments, where $p$ and $N$ are coprime. An intuitive case is for $N=2$, in which the symmetry operation switches fermion parity between even and odd, which is a familiar feature for MZMs.
II.2 Parity conservation
In this subsection, we show a pivotal result that with the generic unitary symmetry protection, as given in Eq. (4), the PZMs can be classified into two nontrivial symmetry sectors after a sophisticated transformation on the parafermion modes, with the parity of each sector being conserved. From Eq. (4) we can show that the unitary symmetry $S$ transforms the PZMs in the way (see Appendix B for more details)
@@ II.2 Parity conservation
where $\mu_{n}^{*}=\mu_{N-n}$, $p_{1}$, $p_{2}$ are integers that satisfy $p_{1}-p_{2}=p$, and $\alpha^{\prime}_{i}=G^{-1}\tilde{\alpha}_{i}G$, $\beta^{\prime}_{i}=G^{-1}\tilde{\beta}_{i}G$. We see that the symmetry operator mixes $\alpha^{\prime}_{i}$ and $\beta^{\prime}_{i}$ in general. We can simply the formula by defining a new set of PZMs via the unitary transformation
$\alpha_{i}=W\alpha^{\prime}_{i}W^{-1},\ \beta_{i}=W\beta^{\prime}_{i}W^{-1},$ (7)
where $W=\exp\bigr{(}{i\sum_{n=1}^{N-1}\nu_{n}Q_{i}^{n}}\bigr{)}$, $\nu_{n}=-\mu_{n}/[(\omega^{n}-1)(\omega^{pn}-1)]$. The generic symmetry protection can be finally rewritten in a decoupled form
$S\alpha_{i}S^{-1}=\omega^{p_{1}}\alpha_{i},\quad S\beta_{i}S^{-1}=\omega^{p_{2}}\beta_{i}.$ (8)
This is an important result that the PZMs can always be written in terms of eigenmodes of the symmetry, even though the parent topological phase hosting such modes is strongly correlatd and cannot be block diagonalized. For the noninteracting phase with MZMs at $N=2$, the above procedure can be easily understood. Namely, for a pair of MZMs $\alpha^{\prime}_{i}$ and $\beta^{\prime}_{i}$, the symmetry acts as $S\alpha^{\prime}_{i}S^{-1}=\pm(\cos\phi\alpha^{\prime}_{i}+\sin\phi\beta^{\prime}_{i})$, $S\beta^{\prime}_{i}S^{-1}=\mp(-\sin\phi\alpha^{\prime}_{i}+\cos\phi\beta^{\prime}_{i})$, which give $\mu_{1}=\phi$. Setting $\nu_{1}=-\phi/4$, we have $\alpha_{i}=\cos\frac{\phi}{2}\alpha^{\prime}_{i}+\sin\frac{\phi}{2}\beta^{\prime}_{i}$, $\beta_{i}=-\sin\frac{\phi}{2}\alpha^{\prime}_{i}+\cos\frac{\phi}{2}\beta^{\prime}_{i}$, and it’s easy to verify that $S\alpha_{i}S^{-1}=\pm\alpha_{i}$, $S\beta_{i}S^{-1}=\mp\beta_{i}$. Nevertheless, for the parafermion modes, the generic unitary symmetry transformation $W$ given above is highly nonlinear.
The decoupled form of unitary symmetry protection described in Eq. (8) leads to an important consequence that the parity of each symmetry sector, defined by $Q_{\alpha}=\omega^{(N+1)/2}\alpha_{1}\alpha_{2}^{\dagger}$ and $Q_{\beta}=\omega^{(N+1)/2}\beta_{1}\beta_{2}^{\dagger}$, should be conserved. To see this, we consider the generic non-local interacting terms allowed in the PZMs that
$H_{I}=\sum_{k,l,m,n}\xi_{k,l,m,n}\alpha_{1}^{k}\alpha_{2}^{l}\beta_{1}^{m}\beta_{2}^{n}+h.c.,$ (9)
where $k+l+m+n=0\text{ mod }N$ due to the total parity conservation. The symmetry protection condition in Eq. (8) further requires that $k+l=m+n=0\text{ mod }N$. Then the interaction term reads $H_{I}=\sum_{k,m}\xi_{k,m}Q_{\alpha}^{k}Q_{\beta}^{m}+h.c.$, which commutes with the parity operator of each sector. Note that the interaction term couples $\alpha_{i}$-sector and $\beta_{i}$-sector when $k\neq 0$, $m\neq 0$, and the parity conservation of each sector does not imply the decomposition of the whole interacting system into decoupled subsystems.